# Modern Neuromorphic AI: From Intra-Token to Inter-Token Processing

### v1

@@ VII-A Backpropagation-Based Learning via Teacher-Forcing
In a conventional supervised learning setting, the model is given examples of the form of a sequence of inputs $\bm{x}^{T}=\{\bm{x}[t]\}_{t=1}^{T}$. As shown in Figure 1, these may be obtained by tokenizing and embedding text or other types of data. Since the desired outputs $\bm{x}^{T}$ are given, this type of setting is also known as teacher forcing, as the “teacher” (the data) “forces” the desired behavior on the model [williams1989learning].
@@
Denote the resulting training loss function as $\mathcal{L}(\theta,\bm{x}^{T})$, where $\theta$ collects all the trainable parameters of the model (see (8)). Evaluating the gradient of this loss, $\nabla_{\theta}\mathcal{L}(\theta,\bm{x}^{T})$, requires backpropagating through the model’s computational graph (see, e.g., [simeone2022machine]). In general, NPEs – and thus also NPUs – are recurrent models implementing dynamics of the form (9). Therefore, evaluating the gradient $\nabla_{\theta}\mathcal{L}(\theta,\bm{x}^{T})$ calls for the use of backpropagation through time (BPTT). The forward pass of BPTT goes sequentially through all time steps t=1,2,,Tt=1,2,...,T; and the backward pass follows the opposite order. This sequential processing may prevent the efficient training of models that operate on many tokens $T$ (apart from suffering from issues such as vanishing gradients).
@@
The total number of computations across all parallel processors can be further reduced from order $\mathcal{O}(T^{2})$ to $\mathcal{O}(T)$ via the parallel scan algorithm, which is tailored to the partial-prefix sum form of the dynamics (16) [blelloch1990prefix, gu2023mamba]. This algorithm has the added advantage that it can be also applied when the gains $a[t]$ and $b[t]$ are not stationary due to implementation of gating (see Section V-B). However, in some cases, it may be less preferable for hardware deployments that can benefit from well-optimized kernels for matrix multiplication.
As discussed in Section V-C, due to the presence of the reset mechanism, the LIF-based dynamics (9) cannot be expressed as an SSM, calling for a sequential evaluation of the gradient via BPTT. However, having removed the reset mechanism, the LIF dynamics can be expressed as a convolution, since the LIF model (9) with matrix $\bm{W}_{\text{res}}=\bm{0}$ is linear in the state variables $\bm{u}[t]$ as for SSMs. Therefore, the evaluation of the membrane potential dynamics can be implemented in parallel through convolutions or the parallel scan algorithm as discussed above [fang2023parallel, bal2024pspikessm].
Retaining the reset mechanism while enabling parallelization requires approximation techniques. For example, one can train a neural network that predicts the spike sequence and then use the predictions in lieu of the true spikes for parallel computation [du2024spikingssms].
@@ VII-B Gradient Evaluation via Surrogate Gradients
To obviate this problem, one approach is the straight-through estimator, which replaces the threshold functions with the identity function during the backward pass [bengio2013estimating]. A potentially more effective approach is to use as a replacement a smooth function that serves as a closer approximation, or surrogate, of the original threshold function [neftci2019surrogate]. For example, for the case of binary outputs (4), common choices of approximating functions include the sigmoid function, the arctangent function, and piece-wise linear approximations [bellec2018long, zenke2018superspike].
@@ VII-C Local Learning via Feedback-Based Local Updates
Backpropagation may be prohibitively complex and incompatible with some neuromorphic computing platforms [davies2021advancing]. Furthermore, it is not applicable to control settings in which data cannot “force” desired behavior. In fact, in such scenarios, the model takes actions and is only then rewarded or penalized [fremaux2013reinforcement]. For all these reasons, it is important to consider alternative learning paradigms based on feedback and local adaptation, rather than on backpropagation and teacher forcing objectives. Most of the work on this topic has focused so far on standard SNN models defined by networks of spiking neurons (see Section III-A), and we limit the discussion here to such settings.
The most fundamental local learning rule is spike-timing-Dependent plasticity (STDP), which adjusts weights connecting spiking neurons based on the relative timing of pre- and post-synaptic spikes. STDP is inspired by neuroscientific principles, and operates in a fully unsupervised way [dan2004spike].
However, STDP alone is insufficient for supervised, or goal-directed, learning. Three-factor learning rules extend STDP by incorporating a global reward or error signal, enabling learning from outcomes [gerstner2018eligibility]. Such rules can be derived from first principles for spiking models with probabilistic binary spiking mechanisms as in (5[jang2019introduction]. In this type of models, neurons respond probabilistically to their inputs. The loss function evaluated at the last layer serves as a reward signal to encourage or discourage the realized behavior on the basis of adherence to labeled data or control goals.

### v2

@@ VII-A Backpropagation-Based Learning via Teacher-Forcing
In a conventional supervised learning setting, the model is given examples of the form of a sequence of inputs $\bm{x}^{T}=\{\bm{x}[t]\}_{t=1}^{T}$. As shown in Figure 1, these may be obtained by tokenizing and embedding text or other types of data. Since the desired outputs $\bm{x}^{T}$ are given, this type of setting is also known as teacher forcing, as the “teacher” (the data) “forces” the desired behavior on the model [59].
@@
Denote the resulting training loss function as $\mathcal{L}(\theta,\bm{x}^{T})$, where $\theta$ collects all the trainable parameters of the model (see (8)). Evaluating the gradient of this loss, $\nabla_{\theta}\mathcal{L}(\theta,\bm{x}^{T})$, requires backpropagating through the model’s computational graph (see, e.g., [53]). In general, NPEs – and thus also NPUs – are recurrent models implementing dynamics of the form (9). Therefore, evaluating the gradient $\nabla_{\theta}\mathcal{L}(\theta,\bm{x}^{T})$ calls for the use of backpropagation through time (BPTT). The forward pass of BPTT goes sequentially through all time steps t=1,2,,Tt=1,2,...,T; and the backward pass follows the opposite order. This sequential processing may prevent the efficient training of models that operate on many tokens $T$ (apart from suffering from issues such as vanishing gradients).
@@
The total number of computations across all parallel processors can be further reduced from order $\mathcal{O}(T^{2})$ to $\mathcal{O}(T)$ via the parallel scan algorithm, which is tailored to the partial-prefix sum form of the dynamics (16) [11, 25]. This algorithm has the added advantage that it can be also applied when the gains $a[t]$ and $b[t]$ are not stationary due to implementation of gating (see Section V-B). However, in some cases, it may be less preferable for hardware deployments that can benefit from well-optimized kernels for matrix multiplication.
As discussed in Section V-C, due to the presence of the reset mechanism, the LIF-based dynamics (9) cannot be expressed as an SSM, calling for a sequential evaluation of the gradient via BPTT. However, having removed the reset mechanism, the LIF dynamics can be expressed as a convolution, since the LIF model (9) with matrix $\bm{W}_{\text{res}}=\bm{0}$ is linear in the state variables $\bm{u}[t]$ as for SSMs. Therefore, the evaluation of the membrane potential dynamics can be implemented in parallel through convolutions or the parallel scan algorithm as discussed above [21, 5].
Retaining the reset mechanism while enabling parallelization requires approximation techniques. For example, one can train a neural network that predicts the spike sequence and then use the predictions in lieu of the true spikes for parallel computation [19]. An alternative iterative mechanism to obtain the spiking times was discussed in [65]. An event-driven parallel scan approach is studied in [58].
@@ VII-B Gradient Evaluation via Surrogate Gradients
To obviate this problem, one approach is the straight-through estimator, which replaces the threshold functions with the identity function during the backward pass [9]. A potentially more effective approach is to use as a replacement a smooth function that serves as a closer approximation, or surrogate, of the original threshold function [41]. For example, for the case of binary outputs (4), common choices of approximating functions include the sigmoid function, the arctangent function, and piece-wise linear approximations [8, 64].
@@ VII-C Local Learning via Feedback-Based Local Updates
Backpropagation may be prohibitively complex and incompatible with some neuromorphic computing platforms [17]. Furthermore, it is not applicable to control settings in which data cannot “force” desired behavior. In fact, in such scenarios, the model takes actions and is only then rewarded or penalized [22]. For all these reasons, it is important to consider alternative learning paradigms based on feedback and local adaptation, rather than on backpropagation and teacher forcing objectives. Most of the work on this topic has focused so far on standard SNN models defined by networks of spiking neurons (see Section III-A), and we limit the discussion here to such settings.
The most fundamental local learning rule is spike-timing-Dependent plasticity (STDP), which adjusts weights connecting spiking neurons based on the relative timing of pre- and post-synaptic spikes. STDP is inspired by neuroscientific principles, and operates in a fully unsupervised way [16].
However, STDP alone is insufficient for supervised, or goal-directed, learning. Three-factor learning rules extend STDP by incorporating a global reward or error signal, enabling learning from outcomes [23]. Such rules can be derived from first principles for spiking models with probabilistic binary spiking mechanisms as in (5[34]. In this type of models, neurons respond probabilistically to their inputs. The loss function evaluated at the last layer serves as a reward signal to encourage or discourage the realized behavior on the basis of adherence to labeled data or control goals.